OpenStudy (anonymous):

PROBLEM: A canvas tent is to be sewn to the dimensions as shown. The tent consists of two side panels, each 10 feet by 10 feet; two end panels that are triangles, each 8 feet high and 12 feet at the base; and a floor that is 12 feet by 10 feet. About how many square feet of canvas are used to make the tent (sides, ends, and floor)?

OpenStudy (whpalmer4):

You need to figure out the area of each of the panels, and add them up. The side and floor panels are rectangles, with the area given by A = l * w. The end panels are triangles, with the area given by A = 1/2 b * h. Plug and chug! :-)

OpenStudy (anonymous):

That makes no sense to me, and I have to pass this

OpenStudy (whpalmer4):

 Those are all of the pieces, labeled with their dimensions. I gave you the formulas for finding their areas. Rectangle (which includes squares): A = l * w Triangle: A = 1/2 * b * h Now does it make sense?